OpenStudy (anonymous):

INTRO PHYSICS: Two planets are 3 x 10^7 km apart. Planet A has a mass of 8 x 10^24 kg. Planet B has a mass of 1 x 10^25 kg. What is the gravitational force that Planet A exerts on Planet B?

OpenStudy (aaronq):

\(F=G\dfrac{m_1m_2}{r^2}\) where \(G=6.674×10^{−11} N m^2 /kg^2\)

OpenStudy (aaronq):

you need to convert to SI units first, by the way.

OpenStudy (anonymous):

but would Planet B exert the same force? I'm only looking for the force that Planet A exerts on Planet B because the next question is how much force does Planet B exert on Planet A @aaronq

OpenStudy (aaronq):

it's the same force

OpenStudy (anonymous):

thank you!

OpenStudy (aaronq):

no problem!